OpenStudy (anonymous):

Show all steps in solving the following equation: 4x - 5 = -21 Explain, including first ...., second..., then show that your solution works! This is worth 3 points.

OpenStudy (anonymous):

4x-5=-21 4x=-21+5 4x=-16 x=-16/4 x=-4